OpenStudy (anonymous):

Suppose the population of trout in a certain stretch of a river is 4000. In about how many years will the population of trout be 700 if the decay rate is 35%? Use the equation 700 = (4000)(0.65)x and round the value of x to the nearest year.

OpenStudy (anonymous):

4 years

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

(4000)(0.65)^4

OpenStudy (anonymous):

@arindameducationusc is that right? ^^^

OpenStudy (anonymous):

okay @nabiladh thank you!

OpenStudy (anonymous):

internets lagging so having a hard time seeing responses

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

actually its 1 year 700=4000(0.65)^x 7000/4000=4000/4000(0.65)^x 7/4=(0.65)^x log(7/4)=log(0.65)^x 0.24=xlog(0.65) x=0.24/(0.187) x=1.28 x=1 year

OpenStudy (anonymous):

stuck between the 2 idk good luck

OpenStudy (arindameducationusc):

@lily1021 i am pretty sure ..thatz not the answer.

OpenStudy (anonymous):

hm okay so not 4

OpenStudy (anonymous):

@zzr0ck3r can you help with this?

OpenStudy (zzr0ck3r):

\(700 = (4000)(0.65)^x\\\dfrac{700}{4000}=0.65^x\\\ln(\dfrac{700}{4000})=\ln(0.65)^x=x\ln(0.65)\\ x=\dfrac{\ln(\dfrac{700}{4000})}{\ln(0.65)}\approx 4.04605\) or \(700 = (4000)(0.65)^x\\\dfrac{700}{4000}=0.65^x\\\log_{0.65}(\dfrac{700}{4000})=\log_{0.65}(0.65^x)=x\\x\approx 4.04605\)

OpenStudy (anonymous):

so it is 4 years then

OpenStudy (zzr0ck3r):

Are you asking me if the closest whole number to \(4.04605\) is \(4\)?

OpenStudy (anonymous):

well I'm just confused now because it was 4 then it wasn't then it was 1 then I tried it on my own and got 3 somehow. It really was a rhetorical question.

OpenStudy (zzr0ck3r):

you can check the answer by plugging it in...

OpenStudy (zzr0ck3r):

its always been 4

OpenStudy (anonymous):

But that was really unnecessary..

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

It's not that funny. It was rude.

OpenStudy (zzr0ck3r):

they are not mutually exclusive

OpenStudy (zzr0ck3r):

It was meant to be funny. Sorry if you do not like my humor, focus on the math.

OpenStudy (anonymous):

Meaning and intent aren't the same thing, so oop @ you. I'm not bothered though.

OpenStudy (zzr0ck3r):

Meaning is actually synonymous to intent?

OpenStudy (arindameducationusc):

yes... @zzr0ck3r steps look good yes thats a confirmation @lily1021

OpenStudy (zzr0ck3r):

no need for confirmation. check your answers! If I am right, then \(4000*0.65^{4.04605}=700\) Get your calculators out!!!

OpenStudy (arindameducationusc):

LOL

OpenStudy (arindameducationusc):

yes, you are right @zzr0ck3r

OpenStudy (anonymous):

Not really how something means to me isn't the same as how you meant something to be. English isn't my first language so maybe that's why you didn't understand.

OpenStudy (anonymous):

Just leave my question, it's done with.

OpenStudy (anonymous):

Please leave my question.

OpenStudy (zzr0ck3r):

I fully understand that you are being a baby. Just say thanks for the help and move on :) It was funny. I know you were not asking me to round it for you, but I wanted to point out that it was a weird question. Also your reply about not being sure because it was 1, then it was 4... did you think I would say, oh actually yeah I am wrong and that other dude is right?

OpenStudy (zzr0ck3r):

jokes again^^^

OpenStudy (arindameducationusc):

:D

OpenStudy (arindameducationusc):

k, guys i am going.... :)

OpenStudy (anonymous):

No you are being rude. So stop. You seem to be the one who can't move on. Like why are you still here? This is my question I politely ask you go away.

OpenStudy (zzr0ck3r):

I can not tell you anymore that I am joking. Also, the question is done. I solved it!!!

OpenStudy (anonymous):

And I cannot tell you anymore it wasn't funny. So get off my question if the question is done!!

OpenStudy (zzr0ck3r):

gn sorry :)

OpenStudy (zzr0ck3r):

And I assure you many people would think it was funny.

OpenStudy (anonymous):

Maybe pretentious douche bags who look down on people who are bad at math.

OpenStudy (anonymous):

Please leave.

OpenStudy (zzr0ck3r):

medal please?

OpenStudy (anonymous):

lol

OpenStudy (zzr0ck3r):

there ya go:)